# On Giant’s Shoulders: Effortless Weak to Strong by Dynamic Logits Fusion

### v1

@@ Abstract
Efficient fine-tuning of large language models for task-specific applications is imperative, yet the vast number of parameters in these models makes their training increasingly challenging. Despite numerous proposals for effective methods, a substantial memory overhead remains for gradient computations during updates. Can we fine-tune a series of task-specific small models and transfer their knowledge directly to a much larger model without additional training? In this paper, we explore weak-to-strong specialization using logit arithmetic, facilitating a direct answer to this question. Existing weak-to-strong methods often employ a static knowledge transfer ratio and a single small model for transferring complex knowledge, which leads to suboptimal performance. To surmount these limitations, we propose a dynamic logit fusion approach that works with a series of task-specific small models, each specialized in a different task. This method adaptively allocates weights among these models at each decoding step, learning the weights through Kullback-Leibler divergence constrained optimization problems. We conduct extensive experiments across various benchmarks in both single-task and multi-task settings, achieving leading results. By transferring expertise from the 7B model to the 13B model, our method closes the performance gap by 96.4% in single-task scenarios and by 86.3% in multi-task scenarios compared to full fine-tuning of the 13B model. Notably, we achieve surpassing performance on unseen tasks. Moreover, we further demonstrate that our method can effortlessly integrate in-context learning for single tasks and task arithmetic for multi-task scenarios.  [^1] Our implementation is available in https://github.com/Facico/Dynamic-Logit-Fusion
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have shown impressive performance in a wide range of tasks [5, 47] including code generation [10, 41], mathematical reasoning [34, 2], tool-use abilities [46, 42], etc. However, training such LLMs requires substantial computational resources, often involving thousands of GPUs and processing trillions of tokens [29, 39], making the adaptation of the base model for new knowledge inefficient. To address these challenges, parameter-efficient tuning methods [15, 22, 11] have emerged, aiming to achieve comparable performance to full fine-tuning while reducing GPU memory requirements. However, challenges persist in tuning and deploying large-scale models on common hardware, as they still involve computation-intensive processes like gradient calculation and back-propagation. Furthermore, these methods may not be feasible when training data are private.
@@ 2.1 Efficient Specialization
Specializing a pretrained model by fine-tuning for downstream tasks has become a primary paradigm [9, 40, 7]. However, with the increasing scale of models, full fine-tuning has become impractical due to the large number of parameters, requiring costly training resources [5, 47].
To tackle this challenge, researchers have developed parameter-efficient tuning methods [27]. These methods aim to achieve performance similar to full fine-tuning while reducing the GPU memory required. They typically involve freezing the original weights and adding task-specific trainable modules [14, 21, 25], low-rank matrices [15, 22, 11], or bias [3]. Despite their benefits, these approaches still demand significant memory for updating gradients and might not be suitable for scenarios where training data are private or inaccessible. Another direction is model merging [49, 16, 51, 35], which trains task-specific models on different domains, and then combines them into a single model at deployment with weight average [50], neuron permutation [1], interpolation [36, 19] or task vectors [16, 52]. However, they typically suffer from parameter interference between different tasks and static optimal solutions may struggle with multiple task domains.
@@ 3.1 Problem Background
where $M^{L}$, $M^{S}$, and $M^{S}_{ft}$ represent the logits of the large model, small model, and fine-tuned small model, respectively. Their normalized distributions are denoted by $P$, $Q$, and $Q_{ft}$, respectively. Here, $\alpha$ is a pre-adjusted hyperparameter that controls the extent of knowledge transfer from the small model. Our analysis from Appendix B demonstrates that logit arithmetic attempts to approximate the shift $(\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})})^{\alpha}$ between the fine-tuned distribution and the pretrained distribution by controlling the parameter $\alpha$ before inference.
@@ 3.3 Extending to the Fusion of Multiple SLMs
where $Q_{t}$ represents the distribution of $\theta_{t}^{S}$ from the $t$-th domain, $\text{Joint}(\{Q_{1...T}\})$ represents joint distribution of $Q_{1},Q_{2},...,Q_{T}$. When we impose constraints like Equation (5), it attempts to align the joint distributions of the logits from the domain-specific small models. However, the distributions of the logits from the domain-specific small models are usually not independent of each other, so their joint distribution cannot be estimated through simple multiplication.
Due to the difficulty in obtaining a joint distribution for multiple expert models, we decompose the joint distribution constraint problem into a multi-object marginal distribution optimization. By aligning the distributions of each domain-specific small model, we can infer an approximate optimal solution within the extension of Equation (2), as shown by:
$\operatorname*{arg\,min}_{\tilde{P}}\sum_{t=1}^{T}\left[KL(\tilde{P}||P)-KL(Q_{t}||Q)\right)^{2}+\left(KL(P||\tilde{P})-KL(Q||Q_{t})\right]^{2}$ (6)
@@
Here we use $M^{S}_{t}$ to represent the logit of $t$-th expert. Our algorithm is outlined in pseudo-code in Algorithm 1 in Appendix C.
@@ 4.1 Datasets
Following the setting in  Liu et al. [32], we evaluate on the following datasets: mathmetical reasoning (GSM8K [8]); factual accuracy (TruthfulQA [31]); realistic knowledge (TriviaQA [20]); multi-domain general knowledge (MMLU benchmark [12]); summarization (CNN-DailyMail (CNN/DM) [43]). All datasets are tested using a 0-shot setting. Detailed information is provided in Appendix D.
@@ 4.3 Baselines
To demonstrate the effectiveness of our method in efficiently transferring knowledge from a small model to a large model, we compare it against a small model fine-tuned in a single domain and a large model without fine-tuning. We also report the performance of a large model fine-tuned on a single domain or fine-tuned using LoRA to demonstrate the feasibility of our method as a replacement for fine-tuning approaches. Additionally, we compare our method with Proxy Tuning (setting $\alpha=1.0$ as in the original work) to highlight the superiority of our method in the transfer process. We use the LLaMA2-chat model as the expert for TruthfulQA (therefore, there is no corresponding LoRA model). For other experts, we use models fine-tuned on the respective training sets. In multi-task scenarios, we follow the common training practice of mixing the training data, with the results serving as our multi-task baseline. More details can be found in the Appendix D.
@@ Appendix A Model Architecture Diagram
Appendix B Proof of Method
@@
$=\ln{P(x_{k}\mid x_{<k})}+\alpha\cdot(\ln{Q_{ft}(x_{k}\mid x_{<k})}-\ln{Q(x_{k}\mid x_{<k})})$
$=\ln{\Biggl{(}P(x_{k}\mid x_{<k})\cdot\Biggl{(}\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})}\Biggr{)}^{\alpha}\Biggr{)}}$
@@
Essentially, it assumes a predefined distribution shift $(\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})})^{\alpha}$ between the fine-tuned and pretrained distributions of a small model, which can construct the trajectory of $\tilde{P}$. It then attempts to approximate the shift for the fine-tuned and pretrained distributions of the large model $\frac{\tilde{P}(x_{k}\mid x_{<k})}{P(x_{k}\mid x_{<k})}$,.
Appendix C Pseudo Code

### v2

@@ Abstract
Efficient fine-tuning of large language models for task-specific applications is imperative, yet the vast number of parameters in these models makes their training increasingly challenging. Despite numerous proposals for effective methods, a substantial memory overhead remains for gradient computations during updates. Can we fine-tune a series of task-specific small models and transfer their knowledge directly to a much larger model without additional training? In this paper, we explore weak-to-strong specialization using logit arithmetic, facilitating a direct answer to this question. Existing weak-to-strong methods often employ a static knowledge transfer ratio and a single small model for transferring complex knowledge, which leads to suboptimal performance. To surmount these limitations, we propose a dynamic logit fusion approach that works with a series of task-specific small models, each specialized in a different task. This method adaptively allocates weights among these models at each decoding step, learning the weights through Kullback-Leibler divergence constrained optimization problems. We conduct extensive experiments across various benchmarks in both single-task and multi-task settings, achieving leading results. By transferring expertise from the 7B model to the 13B model, our method closes the performance gap by 96.4% in single-task scenarios and by 86.3% in multi-task scenarios compared to full fine-tuning of the 13B model. Notably, we achieve surpassing performance on unseen tasks. Moreover, we further demonstrate that our method can effortlessly integrate in-context learning for single tasks and task arithmetic for multi-task scenarios.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have shown impressive performance in a wide range of tasks [5, 54, 50, 52, 36, 51, 61, 62], including code generation [11, 45], mathematical reasoning [37, 2], tool-use abilities [53, 46], etc. However, training such LLMs requires substantial computational resources, often involving thousands of GPUs and processing trillions of tokens [30, 42], making the adaptation of the base model for new knowledge inefficient. To address these challenges, parameter-efficient tuning methods [16, 23, 12] have emerged, aiming to achieve comparable performance to full fine-tuning while reducing GPU memory requirements. However, challenges persist in tuning and deploying large-scale models on common hardware, as they still involve computation-intensive processes like gradient calculation and back-propagation. Furthermore, these methods may not be feasible when training data are private.
@@ 2.1 Efficient Specialization
Specializing a pretrained model by fine-tuning for downstream tasks has become a primary paradigm [9, 44, 7, 10]. However, with the increasing scale of models, full fine-tuning has become impractical due to the large number of parameters, requiring costly training resources [5, 54].
To tackle this challenge, researchers have developed parameter-efficient tuning methods [28]. These methods aim to achieve performance similar to full fine-tuning while reducing the GPU memory required. They typically involve freezing the original weights and adding task-specific trainable modules [15, 22, 26], low-rank matrices [16, 23, 12], or bias [3]. Despite their benefits, these approaches still demand significant memory for updating gradients and might not be suitable for scenarios where training data are private or inaccessible. Another direction is model merging [56, 17, 58, 38, 35], which trains task-specific models on different domains, and then combines them into a single model at deployment with weight average [57], neuron permutation [1], interpolation [39, 20] or task vectors [17, 59, 35]. However, they typically suffer from parameter interference between different tasks and static optimal solutions may struggle with multiple task domains.
@@ 3.1 Problem Background
where $M^{L}$, $M^{S}$, and $M^{S}_{ft}$ represent the logits of the large model, small model, and fine-tuned small model, respectively. Their corresponding normalized distributions are denoted by $P$, $Q$, and $Q_{ft}$. The detailed theoretical proof supporting this logit arithmetic formula is provided in Appendix B. Here, $\alpha$ is a pre-adjusted hyperparameter that controls the extent of knowledge transfer from the small model. Our analysis from Appendix C demonstrates that logit arithmetic attempts to approximate the shift $(\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})})^{\alpha}$ between the fine-tuned distribution and the pretrained distribution by controlling the parameter $\alpha$ before inference.
@@ 3.3 Extending to the Fusion of Multiple SLMs
where $Q_{t}$ represents the distribution of $\theta_{t}^{S}$ from the $t$-th domain, $\text{Joint}(\{Q_{1...T}\})$ represents the distribution of the combined knowledge of $Q_{1},Q_{2},...,Q_{T}$. When we impose constraints like Equation (5), it attempts to align the joint distributions of the logits from the domain-specific small models. However, the distributions of the logits from the domain-specific small models are usually not independent of each other, so their joint distribution cannot be estimated through simple multiplication.
Due to the difficulty in obtaining a joint distribution for multiple expert models, we decompose the joint distribution constraint problem into a multi-object marginal distribution optimization. The transformation process we prove in detail in Appendix E. By aligning the distributions of each domain-specific small model, we can infer an approximate optimal solution within the extension of Equation (2), as shown by:
$\operatorname*{arg\,min}_{\tilde{P}}\sum_{t=1}^{T}\left[\left(D_{KL}(\tilde{P}||P)-D_{KL}(Q_{t}||Q)\right)^{2}+\left(D_{KL}(P||\tilde{P})-D_{KL}(Q||Q_{t})\right)^{2}\right]$ (6)
@@
Here we use $M^{S}_{t}$ to represent the logit of $t$-th expert. Our algorithm is outlined in pseudo-code in Algorithm 1 in Appendix F.
@@ 4.1 Datasets
Following the setting in  Liu et al. [33], we evaluate on the following datasets: mathmetical reasoning (GSM8K [8]); factual accuracy (TruthfulQA [32]); realistic knowledge (TriviaQA [21]); multi-domain general knowledge (MMLU benchmark [13]); summarization (CNN-DailyMail (CNN/DM) [47]). All datasets are tested using a 0-shot setting. Detailed information is provided in Appendix G.
@@ 4.3 Baselines
To demonstrate the effectiveness of our method in efficiently transferring knowledge from a small model to a large model, we compare it against a small model fine-tuned in a single domain and a large model without fine-tuning. We also report the performance of a large model fine-tuned on a single domain or fine-tuned using LoRA to demonstrate the feasibility of our method as a replacement for fine-tuning approaches. Additionally, we compare our method with Proxy Tuning (setting $\alpha=1.0$ as in the original work) to highlight the superiority of our method in the transfer process. We use the LLaMA2-chat model as the expert for TruthfulQA (therefore, there is no corresponding LoRA model). For other experts, we use models fine-tuned on the respective training sets. In multi-task scenarios, we follow the common training practice of mixing the training data, with the results serving as our multi-task baseline. More details can be found in the Appendix G.
@@ 6 Conclusions
7 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62276110, No. 62172039 and in part by the fund of Joint Laboratory of HUST and Pingan Property & Casualty Research (HPL). We thank the Shanghai AI Laboratory for supporting GPU resources. The authors would also like to thank the anonymous reviewers for their comments on improving the quality of this paper.
@@ Appendix A Model Architecture Diagram
Appendix B Proof of Logic Arithmetic
Reinforcement Learning from Human Feedback (RLHF) is commonly employed in optimizing language models, where the learning process involves maximizing a reward objective while penalizing the KL divergence between the learned policy and a reference policy with a coefficient $\beta$:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\mathrm{\theta}}(y\mid x)}[r_{\phi}(x,y)]-\beta D_{\mathrm{KL}}\left[\pi_{\mathrm{\theta}}(y\mid x)\|\pi_{\mathrm{ref}}(y\mid x)\right]$ (7)
Following the DPO [43] framework, we can reformulate the above objective as:
$\begin{split}&\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi_{\mathrm{\theta}}(y\mid x)}\left[r_{\phi}(x,y)\right]-\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi_{\mathrm{\theta}}(y\mid x)}\left[\beta\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]\\ &=\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi_{\mathrm{\theta}}(y\mid x)}\left[r_{\phi}(x,y)-\beta\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]\\ &=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y\mid x)}\left[\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}-\frac{1}{\beta}r_{\phi}(x,y)\right]\\ &=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y\mid x)}\left[\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}-\log\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\right]\\ &=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y\mid x)}\left[\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)}\right]\\ &=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y\mid x)}\left[\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\frac{1}{Z(x)}Z(x)}\right]\\ &=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y\mid x)}\left[\log\frac{\pi(y\mid x)}{\frac{1}{Z(x)}\pi_{\mathrm{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)}-\log Z(x)\right]\end{split}$ (8)
where $Z(x)=\sum_{y}\pi_{\text{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)$ is the partition function. Note that the partition function is a function of only $x$ and the reference policy $\pi_{\text{ref}}$, but does not depend on the policy $\pi_{\theta}$. We can now define
${\pi^{*}(y\mid x)}={\frac{1}{Z(x)}\pi_{\text{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)},$ (9)
Then the equation is convert to
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi(y|x)}\left[\log\frac{\pi(y|x)}{\pi^{*}(y|x)}\right]-\log Z(x)\right]=$ (10)
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[{D_{\mathrm{KL}}(\pi(y|x)\mid\mid{\pi^{*}(y|x)})}-\log Z(x)\right]$
Hence we have the optimal solution:
$\pi(y|x)={\pi^{*}(y|x)}=\frac{1}{Z(x)}\pi_{\text{ref}}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)$ (11)
for all $x\in\mathcal{D}$.
According to Rafailov et al. [43], reward models can be reparameterized as: $r(x,y)=\beta\log\frac{\pi(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ for some model $\pi(y\mid x)$ and a given reference model $\pi_{\text{ref}}(y\mid x)$. Additionally, given a pretrained language model $\pi$ and its fine-tuned counterpart $\pi_{\text{ft}}$, the following relationship holds,
$\pi_{\mathrm{ft}}(y\mid x)=\pi(y\mid x)\exp(\underbrace{\log\frac{\pi_{\mathrm{ft}}(y\mid x)}{\pi(y\mid x)}}_{\text{Implicit reward }})=\pi(y\mid x)\exp(\frac{1}{\beta}r_{\text{ft}}(x,y))$ (12)
This indicates that the fine-tuned model $\pi_{\text{ft}}$ can be interpreted as the solution to constrained reinforcement learning problem with a constraint on the pre-trained model. Consequently, the theoretical framework is applicable to any fine-tuned model, providing an RL-based interpretation of fine-tuning.
We can further introduce an $S$-size implicit reward for $L$-size finetuned models:
$\pi^{L}_{\text{ft}}(y\mid x)$ $=\frac{1}{Z(x)}\pi^{L}(y\mid x)\exp\bigl{(}\frac{1}{\beta}r^{S}_{\text{ft}}(x,y)\bigr{)}$ (13)
$\propto\pi^{L}(y\mid x)\exp\bigl{(}\log\frac{\pi^{S}_{\text{ft}}(y\mid x)}{\pi^{S}(y\mid x)}\bigr{)}$
$\propto\pi^{L}(y\mid x)\frac{\pi^{S}_{\text{ft}}(y\mid x)}{\pi^{S}(y\mid x)}$
By taking the logarithms, we derive the following logit arithmetic formula:
$M^{L}_{\text{ft}}(x)\propto M^{L}(x)+(M^{S}_{\text{ft}}(x)-M^{S}(x))$ (14)
where $M^{L}$, $M^{S}$, and $M^{S}_{ft}$ represent the logits of the large model, small model, and fine-tuned small model, respectively. This completes the derivation of the Equation (2).
Appendix C Proof of Method in Section 3.1
@@ Appendix B Proof of Logic Arithmetic
$\propto\ln{P(x_{k}\mid x_{<k})}+\alpha\cdot(\ln{Q_{ft}(x_{k}\mid x_{<k})}-\ln{Q(x_{k}\mid x_{<k})})$
$\propto\ln{\Biggl{(}P(x_{k}\mid x_{<k})\cdot\Biggl{(}\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})}\Biggr{)}^{\alpha}\Biggr{)}}$
@@
Essentially, it assumes a predefined distribution shift $(\frac{Q_{ft}(x_{k}\mid x_{<k})}{Q(x_{k}\mid x_{<k})})^{\alpha}$ between the fine-tuned and pretrained distributions of a small model, which can construct the trajectory of $\tilde{P}$. It then attempts to approximate the shift for the fine-tuned and pretrained distributions of the large model $\frac{\tilde{P}(x_{k}\mid x_{<k})}{P(x_{k}\mid x_{<k})}$.
Appendix D Efficiency Analysis of the Forward Process
In section 5.2, we conducted an Efficiency Analysis of logit Arithmetic. To better illustrate our efficiency, we further analyze the overall efficiency of our method here.
Overall, during a single forward pass, our method has a similar time complexity to the static method. Given: current sequence length $s$, large model dimension $h_{L}$, small model dimension $h_{S}$, number of layers in the large model $L_{1}$, number of layers in the small model $L_{2}$, batch size $B$, vocabulary size $V$, number of searches per decoding step $n$. Assume the FLOPs for a single forward pass of the large model and the small model are $FLOPs_{L}$ and $FLOPs_{S}$, respectively. The FLOPs can be calculated as: $FLOPs_{L}=L_{1}*(12Bsh_{L}^{2}+2Bs^{2}h_{L})+Bsh_{L}V$ ,and $FLOPs_{S}=L_{2}*(12Bsh_{S}^{2}+2Bs^{2}h_{S})+Bsh_{S}V$(here we ignore the kv cache). Therefore, the FLOPs for a single forward pass of our method on a single task is: $FLOPs_{L}+2*FLOPs_{S}+nBV$. Among these, only the $nBV$ term ($n\leq 20$) corresponds to the additional computational cost of our method, which is much smaller compared to the previous term and can be considered negligible in the overall time. Additionally, in our efficiency analysis, as shown in Table 3, our method is only 0.008 seconds slower per sample compared to the static method, which is negligible.
Appendix E Proof for the Fusion of Multiple SLMs Scenario
This section mainly explains how we extend the transfer problem to multiple small models. When transferring the knowledge of multiple expert SLMs to a LLM, we consider the following two aspects: 1. The fusion of knowledge from different domain experts. 2. The transfer of knowledge from SLM to LLM, i.e., the transfer of knowledge from a single expert, which was discussed in Section 3.2. Intuitively, we first focus on the fusion of different domain experts' knowledge before performing the transfer. Here, we define the distribution of the combined knowledge of these small models as $J$. Therefore, we aim to achieve $D_{KL}(P||\tilde{P})=D_{KL}(Q||J)$.
Since solving for $J$ is difficult, we propose constraining it based on the relationship between $J$ and $\{Q_{i}\}$ to approximate it. Here, we can transform $D_{KL}(Q||J)$ into $D_{KL}(Q||Q_{i})+C_{J}(Q_{i})$, where $C_{J}(Q_{i})$ is the bias function from $Q_{i}$ to $J$. When we approximate $J$ as the centroid of $\{Q_{i}\}$ on the KL-constrained plane, we can implicitly solve these bias functions. According to the definition of the centroid, $J$ can be solved by minimizing the sum of the squared distances to each point, as shown below:
$\arg\min_{J}\sum_{i=1}^{T}(D_{KL}(Q\parallel J)-D_{KL}\left(Q\parallel Q_{i}\right))^{2}\$ (19)
Since our goal is $D_{KL}(P\parallel\tilde{P})=D_{KL}(Q||J)$, substituting this into our equation gives us our final optimization objective:
$\arg\min_{\tilde{P}}\sum_{i=1}^{T}(D_{KL}(P\parallel\tilde{P})-D_{KL}\left(Q_{i}\parallel Q\right))^{2}\$ (20)
To prove the reasonableness of our approximation, we provide a more rigorous proof below. Our initial objective is as follows:
$\arg\min_{\tilde{P}}\sum_{i=1}^{T}(D_{KL}(\tilde{P}\parallel P)-D_{KL}(J||Q))^{2}$ (21)
By assuming $D_{KL}(Q||J)=D_{KL}(Q||Q_{i})+C_{J}(Q_{i})$, we can transform the original problem
$\arg\min_{\tilde{P}}(D_{KL}(\tilde{P}\parallel P)-D_{KL}(J||Q))^{2}$ (22)
into $T$ constrained optimization problems:
$\arg\min_{\tilde{P}}(D_{KL}(\tilde{P}\parallel P)-D_{KL}\left(Q_{i}\parallel Q\right)-C_{J}(Q_{1}))^{2}$ (23)
$\arg\min_{\tilde{P}}(D_{KL}(\tilde{P}\parallel P)-D_{KL}\left(Q_{i}\parallel Q\right)-C_{J}(Q_{T}))^{2}$
After jointly optimizing them, we have:
$\arg\min_{\tilde{P}}\sum_{i=1}^{T}(D_{KL}(\tilde{P}\parallel P)-D_{KL}\left(Q_{i}\parallel Q\right)-C_{J}(Q_{i}))^{2}$ (24)
$\sum_{i=1}^{T}(D_{KL}(\tilde{P}\parallel P)-D_{KL}\left(Q_{i}\parallel Q\right)-C_{J}(Q_{i}))^{2}$
$\leq\sum_{i=1}^{T}(D_{KL}(\tilde{P}\parallel P)-D_{KL}\left(Q_{i}\parallel Q\right))^{2}+\sum_{i=1}^{T}C_{J}(Q_{i}))^{2}$
$=\sum_{i=1}^{T}(D_{KL}(\tilde{P}\parallel P)-KL\left(Q_{i}\parallel Q\right))^{2}+C_{J-Q}$
Since $C_{J-Q}$ is a constant term independent of $\tilde{P}$, we can ignore it. Finally, we solve the original problem by optimizing this upper bound. When we symmetrize the terms in the KL divergence, we can obtain a similar conclusion. Therefore, in the multi-task setting, we can solve it using the following formula: (As shown in Equation (6)):
$\arg\min_{\tilde{P}}\sum_{i=1}^{T}\left[(KL(P\parallel\tilde{P})-D_{KL}\left(Q_{i}\parallel Q\right))^{2}+(KL(\tilde{P}\parallel P)-KL\left(Q\parallel Q_{i}\right))^{2}\right]\$ (25)
Appendix F Pseudo Code